A conical container has a height of 9 cm and a diameter of 6 cm and is filled to the brim with lemonade. It is leading lemonade at a rate of 1 cubic centimer per minute. What is the height of the remaining lemonage after 1 hour.
Answer by mananth(16946)

Volume V = 1/3 * pi*r^2 * h
r=3cm
h=9cm
V= 1/3 * pi*3^2*9
V= 84.82 cm^3
after 1 hour it has drained 60cu.cm
..
84.82-60=24.82 cu.cm is remaining